If f is one-to-one, find an equation for its inverse. f(x)=5x^3+6
Thanks!

Answer by stanbon(75887)

If f is one-to-one, find an equation for its inverse. f(x)=5x^3+6
-----------------
f is one-to-one.
--------------------
Find the Inverse:
1st: Interchange x and y:
x = 5y^3+6
-----
2nd: Solve for "y":
5y^3 = x-6
y^3 = (x-6)/5
----
y = cbrt[(x-6)/5] is the inverse.
======================================
Cheers,
Stan H.
---------------------